OpenStudy (anonymous):

why doesnt 6(2y+6)=4(9+3y) has no solution?

Nnesha (nnesha):

first you have to distribute brackets

Nnesha (nnesha):

when you distribute both brackets left and right then you can see that y cancel out

Nnesha (nnesha):

@salma1234

OpenStudy (anonymous):

You end up having 12y + 36 = 36 + 12y Because you have the exact same thing on both sides, that means you have infinite solutions.

OpenStudy (shadowlegendx):

12y + 36 = 36 + 12y Infinite Solutions, as they are the same on both sides

Nnesha (nnesha):

 you can solve more if you want to